OpenStudy (anonymous):

An electric oven has a resistance of 101 Ω and a voltage of 220 V. How much current does it draw? 2.2 A , 22 A , 220 A , or 2200 A do I just do 220/101 = 2.178.. about 2.2 so 2.2 A is the solution?

OpenStudy (michele_laino):

yes! here we have to apply the Ohm's law, so we have to do this computation: I=V/R= 220/101=...amperes

OpenStudy (anonymous):

ooh yay!! so 2.2! thank you1!:)

OpenStudy (michele_laino):

yes!